How to do with TI-83 Plus?

In class we examined the linear relationship between shoe size and height for females. When looking at the male data, the following statistics were computed. For shoe size, the sample mean was 11 and the sample standard deviation was 1.35. For height, the sample mean was 71.5 and the sample standard deviation was 2.8. The sample correlation between shoe size and height was found to be r = 0.64. Find the regression equation for predicting height from shoe size, and use it to compute predicted values of height.

Compute the slope of the regression line. Round your answer to 3 decimal places.
Answer #1

Given that,

For shoe size (X) : sample mean (\bar x) = 11 and sample standard deviation (Sx) = 1.35

For Height (Y) : sample mean (\bar y) = 71.5 and sample standard deviation (Sy) = 2.8

The sample correlation between shoe size and height was found to be r = 0.64.

We want to find the regression equation for predicting height from shoe size, and use it to compute predicted values of height.

The slope (b1) of the regression equation is,

2.8 bı = r * = 0.64 * 130 = 1.327407 => b1 = 1.327 (rounded to three decimal places)

=> Slope = 1.327

Note : In TI-83 plus calculator we just have to enter values and press ENTER :

0.64 * 2.8/1.35 = 1.327407407

TEXAS INSTRUMENTS TI-83 Plus 0.64*2 STAT PLOT F1 TBLSET F2 FORMAT F3 CALC F4 Y- WINDOW ZOOM TRACE TABLE F5 GRAPH QUIT INS 2nd